OpenStudy (anonymous):

A person can see the top of a building at an angle of 65°. The person is standing 50 ft away from the building and has an eye level of 5 ft. How tall is the building to the nearest tenth of a foot?

OpenStudy (anonymous):

tan65 = y/ 50 -> y = 107 ft 107 + 5 = 112 ft